OpenStudy (anonymous):

can antone tell me if i am right on these please ??? Simplify : ( -3xy4)( -4x3y6) answer 12x^4y^10 Simplify (7x3y2)2 Answer 49x^6y^4 Write 243 in exponential notation. A) 243 = 2.43 x 10^2

OpenStudy (precal):

yes on number 1

OpenStudy (precal):

yes on number 2

OpenStudy (precal):

your 3rd one is scientific notation but it is correct

OpenStudy (anonymous):

thank you i was scared i did them wrong

OpenStudy (precal):

no you did them correct. Remember you are using the laws of exponents

OpenStudy (anonymous):

5)is this one right ? Simplify [ ( 1/4)2 – ( 1/2 )3 ] x 32 A) -1 + 32 = - 2 16 1

OpenStudy (anonymous):

the 16 and the one is suppose to be under the -1 and the 32

OpenStudy (phi):

is this the problem: \[ ( (\frac{1}{4})^2 - (\frac{1}{2})^3) \times 32 \] If so , the (1/4)^2 means (1/4)*(1/4) = 1/16 and (1/2)^3 means (1/2)(1/2)(1/2)= 1/8 = 2/16 so the problem is \[ ( \frac{1}{16} - \frac{2}{16}) \times 32 = \frac{1-2}{16}\times 32= -1\times2= -2\]